OpenStudy (anonymous):

Use the quadratic formula to solve for x: x+1/x=3

OpenStudy (anonymous):

Simplify the left side by finding LCM.

OpenStudy (anonymous):

You'll get (x^2 + 1)/x = 3 x^2 + 1 = 3x x^2 - 3x + 1 = 0 Now use quadratic formula.

OpenStudy (anonymous):

oh ok, thanks =)

OpenStudy (anonymous):

but where did you get x^2 from? it was only x.. o.O

jhonyy9 (jhonyy9):

but you need to push the restriction first of all becauese you work with fractions - yes ?

jhonyy9 (jhonyy9):

so what value not can getting never x and why ?

OpenStudy (anonymous):

We work with quadratic equations.. but there was only x, but somehow it has to become x^2, but I don't really know how...

jhonyy9 (jhonyy9):

how is right (x+1)/x=3 --- hence correct ?

OpenStudy (anonymous):

the problem was just written like this: 

jhonyy9 (jhonyy9):

so if this is right than how you need to make these calcules - first multiply both sides by x ,than will get x(x+1) ------ = 3x --- now simplifie x from numerator with x on denominator x than will get x+1=3x --- now subtract from both sides x x+1 -x=3x-x 1=2x --- so now divide both sides by 2 1/2 =x so this will be the result

OpenStudy (anonymous):

ok thanks :)